# Multiple Imputation of Hierarchical Nonlinear Time Series Data with an Application to School Enrollment Data

### v1

@@ 6 Discussion
We developed a multiple imputation method for hierarchical time series data that can accommodate a specific type of nonlinear relationship between variables and evaluated the performance of the proposed method using a simulation study and an application to a data set on secondary school enrollment rates. Through comparisons with existing methods for multiple imputation of hierarchical time series data, we found that the proposed MINTS method can lead to substantial gains in performance when variables in the imputation model have a nonlinear relationship that can be well-approximated by a piecewise linear function.
@@
In the simulation study, we found that MINTS performed better for estimation of parameters in uncongenial analysis models using nonlinear simulated data compared to existing models based on the MICE and Amelia methodologies. MINTS generally resulted in the smallest MAE and FMI across experiments, while retaining close to nominal coverage in all but the experiments at the highest rate of simulated missingness. The simulation study results suggest that when there is a nonlinear relationship between variables, MINTS is a preferable method over the MICE and Amelia methods. We also found MINTS still performed well when the variables in the imputation model have a linear relationship, but the difference in performance between MINTS and the existing methods was less pronounced.
@@
For the application to the enrollment data, we evaluated how well the MINTS method performs for estimation of parameters in uncongenial analysis models and for prediction of individual missing values. Based on the analysis model validation exercises, we found that MINTS resulted in improved performance for estimation of the regression coefficient in the linear regression of TFR on NER compared to existing models based on the MICE and Amelia methodologies, with the smallest or second smallest MAE in all experiments and close to nominal coverage in all but one experiment. MINTS also had good performance for estimation of the random intercepts regression of TFR on NER at the 10% rate of simulated missingness. However at higher rates of missingness we found that the pan Random Effects method performed better for estimation of the fixed effect coefficient. At all rates of missingness, MINTS tended to have the smallest MAE for point estimation of the variance of the random intercepts. This suggests that for random intercept models, MINTS is likely to be a better choice than the existing methods if a components of variance analysis is of interest, but the pan Random Effects model could be a better choice if only the fixed effect coefficient is of interest. In the out-of-sample validation exercise for the enrollment data, we evaluated how well each imputation method predicted individual missing values for NER. We found that MINTS resulted in substantial improvements in performance compared to the existing methods, with smaller MAE, narrower intervals, and smaller interval scores across experiments. Despite the narrower intervals, MINTS still had close to nominal coverage in all experiments except MNAR 40%. Overall, our results suggest that MINTS is able to balance good predictive performance for individual missing values with good performance for multiply-imputed estimates of parameters in substantive analysis models.
@@
To facilitate easier comparison across validation exercises, we summarize the average MAE across experiments for each multiple imputation method within each validation exercise in Table 4. A version of Table 4 that includes all validation exercises conducted is available in the Appendix. We note that Table 4 should not be interpreted as inferential and is only of interest as an exploratory comparison tool to enable a simple comparison of one metric from the full evaluation results. MINTS consistently results in the smallest average MAE across experiments in each validation exercise. Out of the previously existing imputation methods, we found the models using the pan methodology tend to have the smallest average MAE. For the validation exercises using enrollment data, the pan Random Effects method tends to perform the best out of the previously existing methods, with the second smallest average MAE for the out-of-sample validation exercise and the second or third smallest average MAE for the analysis model validation estimation of $\beta_{1}$. The pan Fixed Effects method also performs well in terms of average MAE for analysis model validation with the enrollment data, but has much larger average MAE for the out-of-sample validation exercise. For the validation exercises using the nonlinear simulated data, the pan Random Effects method has the smallest average MAE out of the existing imputation methods for estimation of $\omega_{1}$.
@@
Method Enrollment Data Nonlinear Data OOS $\beta_{1}$ $\omega_{1}$ MICE PMM 14.68 2.97 20.16 pan Fixed 13.90 0.52 23.18 pan Random 5.36 1.48 14.30 Amelia TS 15.70 2.59 22.76 Amelia CS 9.89 2.15 37.55 Amelia TSCS 7.49 2.04 17.89 MINTS 2.77 0.18 4.15
@@
We note that all of the imputation methods compared assume that the missing data mechanism is ignorable. We conducted sensitivity analyses to evaluate how robust each imputation model is to one type of violation of the ignorability assumption through the MNAR experiments, where the probability of a country-year being simulated as missing was dependent on the observed value. In the simulation studies, we found that all imputation methods considered were relatively robust to the violation of ignorability when the rate of simulated missingness was low. However, substantial increases in bias and undercoverage were seen at the 40% and 80% rates. MINTS tended to be the most robust to these violations for the simulated data but still suffered from bias and undercoverage, especially at the 80% rate. For the enrollment data validation exercises, no imputation method performed consistently well in the MNAR experiments. MINTS and pan Fixed Effects were the most robust to the violation of ignorability for the analysis model validation exercises, but both methods resulted in far below nominal coverage in more than one experiment. In the out-of-sample validation exercise, MINTS and pan Random Effects fared the best in the MNAR experiments. MINTS had the most robust performance in terms of point estimates, while pan Random Effects had the most robust coverage for interval estimates.
@@
In this paper, we focused on the setting where the analysis model is uncongenial to the imputation model. This is a common occurrence for social science data, particularly when the imputation phase is conducted independently from the analysis phase. However, if the analysis of interest is known during the imputation phase, congeniality between analysis and imputation models is a worthwhile goal. Several methods have been proposed for multiple imputation of hierarchical data under congeniality where the imputation model incorporates features of the substantive analysis model, such as Goldstein et al. (2014); Enders et al. (2020); Lüdtke et al. (2020); Grund et al. (2021); among others. Of particular note, Lüdtke et al. (2020) and Grund et al. (2021) propose a method called “mdmb” that uses a sequential decomposition of the joint model. The mdmb method ensures congeniality between analysis and imputation models by incorporating a term representing the substantive analysis model into the sequential decomposition following the substantive-model-compatible philosophy of Bartlett et al. (2015).
MINTS could be extended to congeniality using the same strategy as mdmb by adding a term representing the substantive analysis model into the decomposition of the joint model as
$p(\mathbf{Z},\mathbf{Y},\mathbf{X}|\bm{\theta}_{Z},\bm{\theta})=$ $p(\mathbf{Z}|\mathbf{Y},\mathbf{X},\bm{\theta}_{Z})p(\mathbf{Y}|\mathbf{X},\bm{\theta}_{Y})p(\mathbf{X}|\bm{\theta}_{X}),$
where $\mathbf{Z}$ is the outcome variable for the analysis of interest with parameters $\bm{\theta}_{Z}$. When the exact substantive analysis model of interest is included in the decomposition, MINTS could be adapted to a fully Bayesian approach and parameter estimates for the analysis model could be obtained directly from the estimation phase of the MCMC.
@@ 7 Conclusion
We have proposed the MINTS method for multiple imputation of hierarchical nonlinear time series data in the bivariate setting where one auxiliary variable is used to impute the variable of interest and the variable of interest has a larger amount of missing data than the auxiliary variable. This setting is motivated by a school enrollment rate data set from the World Bank that includes two measures of enrollment. The Net Enrollment Rate (NER) is the variable of interest for an analysis of the relationship between the Total Fertility Rate and NER, but suffers from a large amount of missing data due to the difficulty of measuring NER. A second measure of enrollment, the Gross Enrollment Ratio (GER), is easier to measure and has a smaller amount of missing data. The MINTS method leverages the strong piecewise linear relationship between NER and GER to impute missing values in both NER and GER using a combination of A-splines, country-specific intercepts, and time series methods using a sequential decomposition of the joint model.
@@
We compared MINTS with existing methods for multiple imputation of hierarchical time series data through a simulation study and an application to the enrollment data set. We considered three models within the MICE framework of van Buuren and Groothuis-Oudshoorn (2011) and three models within the Amelia II framework of Honaker and King (2010). We found that MINTS resulted in better performance overall for estimation of parameters in uncongenial linear regression and random intercept models compared to the imputation models based on the MICE and Amelia methodologies in both a simulation study using nonlinear simulated data and the application to school enrollment data, though models based on the pan method of Schafer and Yucel (2002) were strong alternatives. We also conducted an out-of-sample validation exercise for prediction of individual missing values using the enrollment data and found that MINTS resulted in better predictive performance compared to the existing methods. Based on these validation exercises, we believe MINTS has a promising capability to improve the quality of multiply imputed data sets for hierarchical nonlinear time series data.
@@
One limitation of the MINTS method is the use of spline estimation to model the nonlinear relationship between variables in the imputation model. The A-splines method of Goepp et al. (2018) helps to automate the estimation procedure, but as with all spline estimation methodologies there is the risk of overfitting. The MINTS algorithm uses linear splines and modifies the number of starting knots used in the A-spline algorithm when the sample size is small to reduce the risk of overfitting, but the spline estimation settings are likely to require manual tuning for applications of MINTS to other data sets. We note that other curve-fitting methods, such as LOESS, could be used instead.
@@
MINTS also has several practical limitations compared to MICE and Amelia. Currently, MINTS is restricted to the bivariate setting with continuous variables. In principle, MINTS could be extended to the multivariate setting by adding additional univariate conditional terms to the sequential decomposition of the joint distribution, where careful consideration is needed for the ordering of the added conditional distributions. The imputation models for the added variables could be assumed to be linear or could incorporate a separate spline term for each marginal relationship in the conditional imputation model. MINTS could also be extended to accommodate categorical variables through the use of generalized regression models, for example using similar methodology as Lee and Mitra (2016). Another practical limitation of MINTS is computation time, where the estimation of the MINTS model is much more computationally intense than estimation of the MICE models and the two simpler Amelia models. Computation time could be improved by coding the MCMC algorithm in a more efficient language than R. Extending the MINTS methodology to the multivariate and mixed data type setting and improving the computational efficiency of the MINTS sampling algorithm is of interest for future work.
Acknowledgements
@@
Appendix A Simulation of MAR and MNAR
For MAR, we assumed observations in earlier years are more likely to be missing. For each country-time pair $(c,t)$, we simulated a value $A_{c,t}$ based on $t$ with a large noise term. For the nonlinear and linear simulated data, $A_{c,t}\sim N(t,10^{2})$. For the enrollment data, $A_{c,t}\sim N(t,40^{2})$. For rate of simulated missingness $r$, we determined which values to simulate as missing based on the $r$th quantile of $\mathbf{A}$ as follows:

### v2

@@ 6 Discussion
We developed a multiple imputation method for hierarchical time series data that accommodates nonlinear relationships between variables. We evaluated the performance of the proposed method using a simulation study and an application to a data set on secondary school enrollment rates. Through comparisons with existing methods for multiple imputation of hierarchical time series data, we found that the proposed MINTS method can lead to substantial gains in performance when variables in the imputation model have a nonlinear relationship.
@@
In the simulation study, MINTS generally performed better for estimation of parameters in uncongenial analysis models compared to existing models based on the MICE and Amelia methodologies. We found that MINTS is likely to balance the tradeoff between MAE and coverage better than existing methods when there is a nonlinear relationship between variables, though undercoverage is a concern at very high rates of missingness when the number of complete cases is small. MINTS still performed well in validation exercises with a linear relationship between variables, but we found the difference in performance between MINTS and the existing methods was less pronounced.
@@
For the application to the school enrollment data, we evaluated how well MINTS performed for estimation of parameters in uncongenial analysis models and for prediction of individual missing values. MINTS resulted in improved performance for estimation of the linear regression of TFR on NER compared to existing models based on MICE and Amelia, with the smallest or second smallest MAE in all experiments and close to nominal coverage in all but one experiment. MINTS generally had good performance for estimation of the random intercept model regressing TFR on NER, with the smallest MAE for estimation of the variance of the random intercepts across experiments. However, pan Random Effects performed better for estimation of the fixed effect coefficient on NER at higher rates of missingness. For random intercept models, MINTS is likely to be a better choice than existing methods if a components of variance analysis is of interest, but pan Random Effects could be a better choice if only the fixed effect coefficient is of interest. In the out-of-sample validation exercise, we found that MINTS resulted in substantial improvements for prediction of individual missing values of NER compared to existing methods, with smaller MAE, narrower intervals, and smaller interval scores across experiments while maintaining close to nominal coverage in all but one experiment.
@@
To facilitate easier comparisons across validation exercises, we summarize the average MAE across experiments for each multiple imputation model and the data augmentation approach based on MINTS in Table 5. A version of Table 5 that includes all validation exercises conducted is available in Section 8 of the Supplementary Material (Liu and Raftery (2025)). We note that Table 5 should not be interpreted as inferential and is only of interest as an exploratory comparison tool to enable a simple comparison of one metric from the full evaluation results. Out of the multiple imputation methods considered, MINTS consistently results in the smallest average MAE across experiments in each validation exercise. For the validation exercise using the nonlinear simulated data, the data augmentation approach resulted in the smallest average MAE, followed by multiple imputation using MINTS.
@@
Method Enrollment Data Nonlinear Data OOS $\beta_{1}$ $\omega_{1}$ MICE PMM 14.68 2.97 18.28 pan Fixed 13.90 0.52 14.37 pan Random 5.36 1.48 16.15 Amelia TS 15.70 2.59 12.07 Amelia CS 9.89 2.15 29.86 Amelia TSCS 7.49 2.04 12.93 MINTS 2.77 0.18 4.15 MINTS DA 0.38 2.33
@@
We note that all of the multiple imputation methods compared assume that the missing data mechanism is ignorable. We conducted sensitivity analyses to evaluate how robust each imputation model is to one type of violation of the ignorability assumption through the MNAR experiments. In the simulation studies, we found that all imputation methods considered were relatively robust to the violation of ignorability when the rate of simulated missingness was low. However, substantial increases in bias and undercoverage were seen at the 40% and 80% rates. MINTS tended to be the most robust to these violations but still suffered from bias and undercoverage, especially at the 80% rate. No imputation method performed consistently well in the MNAR experiments for the school enrollment data, though we found MINTS and the pan models tended to be the most robust to violations of ignorability. Across validation exercises, the data augmentation approach based on MINTS generally performed better than all of the multiple imputation methods in the MNAR experiments.
Our results suggest that MINTS is generally able to balance good predictive performance for individual missing values with good performance for multiply-imputed estimates of parameters in substantive analysis models. However, MINTS should be used with caution if a strong violation of the MAR assumption is suspected. Additionally, MINTS is not recommended for use if there is a very small number of observed values available to estimate the imputation model. This small sample size setting occurred for the nonlinear and linear simulated data sets, where some replications of the experiments at the 80% rate of simulated missingness had fewer than 15 complete cases of $\mathbf{X}$ and $\mathbf{Y}$ after simulating data as missing. Although our validation exercises suggest MINTS tends to perform better than the existing multiple imputation methods in this small sample size setting, MINTS still resulted in larger bias than the data augmentation approach and had below nominal coverage. If analysts have access to $\mathbf{X}$, $\mathbf{Y}$, and $\mathbf{Z}$, data augmentation could be a better alternative to multiple imputation in settings with a strong violation of the MAR assumption or with a very small number of complete cases. However, the data augmentation approach may not be practical if analysts are interested in more than one analysis using $\mathbf{Y}$. Unlike multiple imputation, the data augmentation approach is specific to a given analysis model and must be run separately for each analysis model of interest.
@@
In this paper, we focused on the setting where the analysis model is uncongenial to the multiple imputation model. This is a common occurrence for social science data, particularly when the imputation phase is conducted independently from the analysis phase. However, if the outcome variable $\mathbf{Z}$ is known during the imputation phase, congeniality between analysis and imputation models is a worthwhile goal. Several methods have been proposed for multiple imputation of hierarchical data under congeniality where the imputation model incorporates features of the substantive analysis model, such as Goldstein, Carpenter and Browne (2014); Enders, Du and Keller (2020); Lüdtke, Robitzsch and West (2020); Grund, Lüdtke and Robitzsch (2021); among others. Many of these methods can account for nonlinear relationships, either through user-specified functional forms (e.g. Erler et al. (2019)) or through automated modeling procedures (e.g. Kim et al. (2014)). Of particular note, Lüdtke, Robitzsch and West (2020) and Grund, Lüdtke and Robitzsch (2021) propose a method called “mdmb” that uses a sequential decomposition of the joint model. The mdmb method ensures congeniality between analysis and imputation models by incorporating a term representing the substantive analysis model into the sequential decomposition following the substantive-model-compatible philosophy of Bartlett et al. (2015). MINTS could be extended to congeniality using the same strategy as mdmb by adding a term representing the substantive analysis model into the decomposition of the joint model as
$p(\mathbf{Z},\mathbf{Y},\mathbf{X}|\bm{\theta}_{Z},\bm{\theta})=$ $p(\mathbf{Z}|\mathbf{Y},\mathbf{X},\bm{\theta}_{Z})p(\mathbf{Y}|\mathbf{X},\bm{\theta}_{Y})p(\mathbf{X}|\bm{\theta}_{X}).$
When $p(\mathbf{Z}|\mathbf{Y},\mathbf{X},\bm{\theta}_{Z})$ is the exact substantive analysis model of interest, this decomposition corresponds to the data augmentation approach considered in our validation exercises.
@@ 7 Conclusion
We have proposed the MINTS method for multiple imputation of hierarchical nonlinear time series data. We considered the bivariate setting where one variable is the variable of interest for analysis that represents a refined measure of an underlying quantity of interest, and the second variable is an auxiliary variable that represents a coarse measure of the same underlying quantity and has a nonlinear relationship with the variable of interest. The refined measure is more difficult to measure and has a greater amount of missing data than the coarse measure. This setting is motivated by a secondary school enrollment data set that includes two measures of school enrollment rates. The Net Enrollment Rate (NER) is the variable of interest for an analysis of the relationship between the Total Fertility Rate and NER. Measurement of NER can be difficult, as NER incorporates information about both the number and ages of children enrolled in school. The Gross Enrollment Ratio (GER) is a second measure school enrollment that only requires knowledge of the number of children enrolled in school. GER is easier to measure and has a smaller amount of missing data than NER. The MINTS method leverages the strong nonlinear relationship between NER and GER to impute missing values in NER using a combination of smoothing splines, country-specific intercepts, and time series methods.
@@
We compared MINTS with several existing methods for multiple imputation of hierarchical time series data through a simulation study and an application to the school enrollment data. We considered three models within the MICE framework of van Buuren and Groothuis-Oudshoorn (2011) and three models within the Amelia II framework of Honaker and King (2010). We found that MINTS generally resulted in better performance for estimation of parameters in uncongenial linear regression and random intercept models compared to imputation models based on the MICE and Amelia methodologies in both a simulation study using nonlinear simulated data and an application to the school enrollment data, though undercoverage can be a concern when the sample size is small. We also conducted an out-of-sample validation exercise for prediction of individual missing values using the school enrollment data and found that MINTS resulted in better predictive performance compared to the existing methods. Based on these validation exercises, we believe MINTS has a promising capability to improve the quality of multiply imputed data sets for hierarchical nonlinear time series data.
@@
One limitation of the MINTS method is the use of spline estimation to model the nonlinear relationship between variables in the imputation model. The A-splines method of Goepp, Bouaziz and Nuel (2018) helps to automate the estimation procedure, but as with all spline estimation methodologies there is the risk of overfitting. The MINTS algorithm uses linear splines and modifies the number of starting knots used in the A-spline algorithm when the sample size is small to reduce the risk of overfitting, but the spline estimation settings are likely to require manual tuning for applications of MINTS to other data sets. The risk of overfitting is especially of concern when the number of complete cases is small. The splines $f$ and $h$ estimated using A-splines should be checked for each application to ensure the estimated relationships make sense based on substantive knowledge of the variables in the imputation model. We note that other curve-fitting methods, such as LOESS, could be used instead. The estimation of $f$ and $h$ in MINTS is also limited by the assumption that the nonlinear relationship between $\mathbf{X}$ and $\mathbf{Y}$ does not change substantially across countries and times. If there are a sufficient number of complete cases in the data set, country- or time-specific estimates of $f$ and $h$ could be used instead.
The illustrations of MINTS presented in the simulation studies and the application to the school enrollment data are additionally limited by the use of a data-based algorithm to specify certain hyperparameters in the prior distributions. Although we found that the resulting prior distributions were sufficiently diffuse and did not overwhelm the posterior for these specific applications, a consequence of using the data-based algorithm is that the posterior distributions used to create multiple imputations are approximate rather than fully Bayesian posterior distributions. Alternative methods for specifying these hyperparameters that do not rely on the data, such as expert elicitation or using information from previous studies, could be used to address this limitation.
@@
MINTS also has several practical limitations compared to MICE and Amelia. Currently, MINTS is restricted to the bivariate setting with continuous variables. While we are primarily motivated by multiple imputation in settings like the school enrollment rates application with two highly related variables that aim to measure the same underlying quantity, incorporation of additional variables into the MINTS model could improve the plausibility of the MAR assumption. In principle, MINTS could be extended to the multivariate setting by adding additional univariate conditional terms to the sequential decomposition of the joint distribution, where careful consideration is needed for the ordering of the added conditional distributions. The imputation models for the added variables could be assumed to be linear or could incorporate a separate spline term for each marginal relationship in the conditional imputation model. MINTS could also be extended to accommodate categorical variables through the use of generalized regression models, for example using similar methodology as Lee and Mitra (2016). Another practical limitation of MINTS is computation time, where the estimation of the MINTS model is much more computationally intense than estimation of the MICE models and the two simpler Amelia models. Computation time could be improved by coding the MCMC algorithm in a more efficient language than R. Extending the MINTS methodology to the multivariate and mixed data type setting and improving the computational efficiency of the MINTS sampling algorithm is of interest for future work.
{acks}
[Acknowledgments] The authors would like to thank the members of the working group on Applied, Bayesian, and Computational Statistics at the University of Washington for their helpful discussion. We also thank the referees, Associate Editor, and Editor for their valuable feedback.
{funding}